OpenStudy (anonymous):

If the rate of change of the function f(t) = t^(4/3) when t=c is twice the rate of change of f(t) when t=8, then c=?

OpenStudy (anonymous):

Does the wording even make sense to you?

OpenStudy (anonymous):

Yes. First find the derivative of f of t. Do you know what that is?

OpenStudy (anonymous):

f'(t)*t' right?

OpenStudy (anonymous):

so, (4/3)t^(1/4)

OpenStudy (anonymous):

exponet is 1/3

OpenStudy (anonymous):

right, typo on my part, thank you.

OpenStudy (anonymous):

Now set up an equation involving c and 8\[\frac{ 4 }{ 3 }(c)^{\frac{ 1 }{ 3 }}=(2)\frac{ 4 }{ 3 }(8)^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

how did you get that equation?

OpenStudy (anonymous):

Since I knew the formula for each rate of change, I substituted c for t on one side and 8 on the other, just as the problem stipulated. And then I added that the rate of change at c is twice that at 8

OpenStudy (anonymous):

Oh, alright, cool

OpenStudy (anonymous):

Can you solve for c then?

OpenStudy (anonymous):

I got 3root4... am I doing something wrong?

OpenStudy (anonymous):

\[\Large f'(t)=\frac{4}{3}t^{\frac{1}{3}}\] when t=c \[\Large f'(t)=\frac{4}{3}c^{\frac{1}{3}}\] now again \[\Large f'(t)=\frac{4}{3}t^{\frac{1}{3}}\] when t=8 then \[\Large f'(t)=\frac{4}{3}8^{\frac{1}{3}}\] since when its t=c its twice so set the two equal \[\Large 2(\frac{4}{3})C^{\frac{1}{3}}=\frac{4}{3}*(8)^{\frac{1}{3}}\] just solve for C

OpenStudy (anonymous):

Sort of, but you're close. Reduce the equation to its bare bones. On one side, the right, you have 2 times 4/3 times the cubic root of 8 (or 2). On the other side, you also have 4/3, which can cancel with the 4/3 on the other side

OpenStudy (anonymous):

I'm trying to hurry here. \[c ^{\frac{ 1 }{ 3 }}=2*2\]

OpenStudy (anonymous):

oh, so 4^3, which is 64

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Last equation in my response is \[\Large \frac{4}{3}C^{\frac{1}{3}}=2 (\frac{4}{3})8^{\frac{1}{3}}\]

OpenStudy (anonymous):

Righty, thank you so much! :D